OpenStudy (rebeccapink6398):

*WILL GIVE MEDAL *MUST SHOW WORK ~multiply and simplify 5 1/4 TIMES 8/9 TIMES 8/21 5 1/4 * 8/9 * 8/21 PS THEY ARE SAME QUESTION

OpenStudy (anonymous):

16/9

OpenStudy (rebeccapink6398):

that's not one of the options. there are the options A. 80/180 B. 1 7/9 C. 5 16/189 D. 5 445/756

OpenStudy (rebeccapink6398):

@bookworm00981 please help

OpenStudy (anonymous):

Oh Nooo! I'm bad at Mathematics!

OpenStudy (rebeccapink6398):

lol

OpenStudy (anonymous):

@linda3 is better at this kind of stuff :P

OpenStudy (rebeccapink6398):

ok then @bjschultze87 and @ryamorgan284 could you guys help plzzz

OpenStudy (anonymous):

b is you answer \[5\frac{ 1 }{ 4 }*\frac{ 8 }{ 9 }*\frac{ 8 }{ 21 }=1\frac{ 588 }{ 756 }\] which then simplifes down to \[1\frac{ 294 }{ 378 }\] which simplifes down to \[1\frac{ 147 }{ 189 }\] again simplies down to \[1\frac{ 49 }{ 63 }\] which will simplify down one last time to \[1\frac{ 7 }{ 9 }\] and that is you answer I hope that helps

OpenStudy (rebeccapink6398):

thank you so much you'r a lifesaver!!! :D

OpenStudy (anonymous):

you're welcome